Describe how you would prepare a 10% solution hydrochloric acid(liquid) and a 10% solution of sodium chloride(crystals) into a total volume of 100cm3.
1
Expert's answer
2019-12-09T11:48:30-0500

Let's have 100 ml of concentrated hydrochloric acid (HCl) solution (W = 36%, density = 1,18 g/ml)

Then,

W(HCl)= x g HCl / 100 ml of solution = x/118 g = 36%

Hence m(HCl) equals x = 0,36•118 = 42,48 g


To prepare 10% solution of HCl we should add to concentrated HCl y grams(ml) of water. Then, we have the proportion:

0,1 = 42,48 g of HCl/118 g of first solution + y grams of H2O

Or 0,1 = 42,48/118+y

y+118=424,8

y = 306,8 g(ml) of H2O

SO, to prepare 10% solution of HCl we should add 306,8 ml of H2O to 100 ml of concentrated HCl. OR, to prepare 100 cm3 of 10% solution of HCl we should add to 23,54 cm3 of concentrated HCl 76,46 cm3 of water.


To prepare 100 ml(cm3) of 10% solution of NaCl:

0,10 = x g NaCl/100 cm3 of solution

x = 10 g

You should take a 10 g of NaCl and 90 g of water.